A loan of $500 is to be paid with two payments of $300, one in 3 months, and another in 6 months. The compound interest charged per-annum on the loan is:
1
Expert's answer
2020-06-21T13:35:03-0400

Loan amount = $500

Amount = 2 payments of 300 = 300*2 = 600

T = 6months = 6/12 = 0.5 years

A = P * (1+r/n)nt

600 = 500 * (1+r)0.5

(1+r)0.5 =600/ 500

0.5 log (1 + r) = log 1.2

Log (1+ r) = log 1.2 /0.5

Log (1+ r) = 0.079181/0.5

1+r = log-10.158362

1+ r = 1.43999836835

r =1.43999836835-1

r = 0.439998 ~ 44% per annum